# ZJUKLAB at SemEval-2025 Task 4: Unlearning via Model Merging

### v1

@@ 2 Task Description
For details about task description, please refer to the official paper (Ramakrishna et al., 2025).
@@ 6.1 Drawbacks: Over-forgetting Phenomena
Despite demonstrating effectiveness, our system still exhibits over-forgetting. Firstly, the unlearned model exhibits model collapse, frequently generating repetitive characters (e.g., "6 6 6"). This phenomenon arises from the training process itself, the model may find a suboptimal but easy shortcut: generating repetitive outputs to reduce loss. Specifically, Task 2 involves a digit-heavy dataset, so the model will take this high-frequency option as their outputs. Secondly, we observed forgetting of generic knowledge. We analyze question patterns of forget set to construct 50 common knowledge questions (e.g., "What is the capital of France?"), finding a significant Knowledge Score drop (0.88 $\to$ 0.35) against a vanilla baseline. These drawbacks are also observed in some studies Mekala et al. (2025); Xu et al. (2025), highlighting a fundamental weakness of this paradigm. By primarily focusing on suppressing the original tokens, the model is constrained from discovering meaningful alternative outputs and consequently falls into a shortcut of generating meaningless responses.
@@ 6.3 Rethinking Unlearning’s Objectives
Recent studies (Xu et al., 2024; Zhou et al., 2024; Thaker et al., 2024; Cooper et al., 2024; Barez et al., 2025) present critical analyses of generative AI unlearning. These studies collectively reveal three fundamental limitations: (1) current unlearning methods remain impractical, (2) evaluations fail to assess the generalization capability of unlearned models, and (3) benchmarks encourage model to overfit the training set, creating an illusory forgetting. The root challenge lies in the absence of a clearly defined and universally applicable unlearning objective. In a pragmatic sense, we should not set too many objectives for unlearning, such as Resilient to Relearning Attacks (Fan et al., 2025). Rather than treating unlearning as an ultimate solution for privacy and copyright issues, future research should develop on-demand unlearning and robust evaluation to address policy concerns.
@@ Acknowledgements
We thank ZJUKLAB for their invaluable support and resources throughout this research project.
@@ Appendix B Related Work
The topic of unlearning in large language models (Chen, 2024) has recently attracted significant attention in the literature. One approach to unlearning is Gradient Ascent (Jang et al., 2023), which aims to maximize the loss function to facilitate forgetting. Another method, Negative Preference Optimization (NPO) (Zhang et al., 2024a), builds upon Direct Preference Optimization (DPO) (Rafailov et al., 2023), offering an alternative strategy for model unlearning. Various unlearning techniques have been proposed, including those presented by (Lu et al., 2022; Eldan and Russinovich, 2023; Yu et al., 2023; Chen and Yang, 2023; Wang et al., 2025; Gandikota et al., 2024; Jiang et al., 2025; Liu et al., 2024b; Zhuang et al., 2024). An alternative strategy, referred to as “locate-then-unlearn,” is exemplified by KnowUnDo (Tian et al., 2024) and SURE (Zhang et al., 2024b), which focus on knowledge localization before executing the unlearning process. Additionally, data-driven methods for unlearning have also been introduced, such as those proposed by (Jang et al., 2022; Ma et al., 2024; Liu et al., 2024a; Gu et al., 2024; Sinha et al., 2024; Xu et al., 2025; Mekala et al., 2025).

### v2

@@ 2 Task Description
For details about task description, please refer to the official paper (Ramakrishna et al., 2025; ramakrishna2025semeval2025task4unlearning).
@@ 6.1 Drawbacks: Over-forgetting Phenomena
Despite demonstrating effectiveness, our system still exhibits over-forgetting. Firstly, the unlearned model exhibits model collapse, frequently generating repetitive characters (e.g., "6 6 6"). This phenomenon arises from the training process itself, the model may find a suboptimal but easy shortcut: generating repetitive outputs to reduce loss. Specifically, Task 2 involves a digit-heavy dataset, so the model will take this high-frequency option as their outputs. Secondly, we observed forgetting of generic knowledge. We analyze question patterns of forget set to construct 50 common knowledge questions (e.g., "What is the capital of France?"), finding a significant Knowledge Score drop (0.88 $\to$ 0.35) against a vanilla baseline. These drawbacks are also observed in some studies Mekala et al. (2025); Xu et al. (2025), highlighting a fundamental weakness of this paradigm. Throughout training, the system repeatedly applies reverse optimization signals to the original forget data. Without positive guidance like in reinforcement learning, the model cannot explore better outputs and inevitably degrades under sustained pressure.
@@ 6.3 Rethinking Unlearning’s Objectives
Recent studies (Xu et al., 2024; Zhou et al., 2024; Thaker et al., 2024; Cooper et al., 2024; Barez et al., 2025) present critical analyses of generative AI unlearning. These studies collectively reveal three fundamental limitations: (1) current unlearning methods remain impractical, (2) evaluations fail to assess the generalization capability of unlearned models, and (3) benchmarks encourage model to overfit the training set, creating an illusory forgetting. The root challenge lies in the lack of a clearly defined, universally applicable unlearning objective. Rather than overloading unlearning with goals like resistance to relearning attacks (Fan et al., 2025), future research should prioritize on-demand unlearning and robust evaluation to address practical policy needs. As discussed in §6.1, current methods often lead to degraded outputs. Future work can explore the incorporation of positive signals to guide the model toward more appropriate forgetting behaviors such as data augmentation and reinforcement learning.
@@ Acknowledgements
We would like to express our great gratitude to the anonymous reviewers for their kind comments. This work was supported by the National Natural Science Foundation of China (No. 62206246), the Fundamental Research Funds for the Central Universities (226-2023-00138), Yongjiang Talent Introduction Programme (2021A-156-G), CIPSC-SMP-Zhipu Large Model Cross-Disciplinary Fund, Tencent AI Lab Rhino-Bird Focused Research Program (RBFR2024003). We gratefully acknowledge the support of Zhejiang University Education Foundation Qizhen Scholar Foundation.
@@ Appendix B Related Work
The topic of unlearning in large language models (Chen, 2024) has recently attracted significant attention in the literature. One approach to unlearning is Gradient Ascent (Jang et al., 2023), which aims to maximize the loss function to facilitate forgetting. Another method, Negative Preference Optimization (NPO) (Zhang et al., 2024a), builds upon Direct Preference Optimization (DPO) (Rafailov et al., 2023), offering an alternative strategy for model unlearning. Various unlearning techniques have been proposed, including those presented by (Lu et al., 2022; Eldan and Russinovich, 2023; Yu et al., 2023; Chen and Yang, 2023; Wang et al., 2025; Gandikota et al., 2024; Jiang et al., 2025; Liu et al., 2024b; Zhuang et al., 2024). An alternative strategy, referred to as “locate-then-unlearn,” is exemplified by KnowUnDo (Tian et al., 2024) and SURE (Zhang et al., 2024b), which focus on knowledge localization before executing the unlearning process. Additionally, data-driven methods for unlearning have also been introduced, such as those proposed by (Jang et al., 2022; Ma et al., 2024; Liu et al., 2024a; Gu et al., 2024; Sinha et al., 2024; Xu et al., 2025; Mekala et al., 2025). Several works have explored the use of model merging techniques to achieve unlearning (kadhe2024splitunlearnmergeleveraging; kuo2025exactunlearningfinetuningdata).